OpenStudy (anonymous):

jackson's property tax is $616.97 and is due April 20th. he does not pay until june 17th. the county adds a penalty of 8.1% simple interest on unpaid tax. find the penalty using exact interest.

jimthompson5910 (jim_thompson5910):

You would use the formula I = Prt to compute the interest

OpenStudy (anonymous):

but what would be t?

jimthompson5910 (jim_thompson5910):

in this case P = 616.97 (principal r = 0.081 (interest) t = 27/365 (this is because you're using exact interest) I used wolfram alpha to count the days from april 20th to june 17th http://www.wolframalpha.com/input/?i=april+20+to+june+17

jimthompson5910 (jim_thompson5910):

oh wait, used the wrong number lol

OpenStudy (anonymous):

oh okay I wasn't sure about that

OpenStudy (anonymous):

right, it should be around 57

jimthompson5910 (jim_thompson5910):

t = 59/365 sorry I misread and the 27 was referring to days from today til april 20th this page has the count you want http://www.timeanddate.com/date/durationresult.html?m1=4&d1=20&y1=&m2=6&d2=17&y2=&ti=on

jimthompson5910 (jim_thompson5910):

So you'll be adding on I = Prt I = 616.97(0.081)*(59/365) I = 8.0780811780822 I = 8.08 dollars in interest to the amount you already owe

OpenStudy (anonymous):

I got it wrong. That's strange.

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

It doesn't say the correct answer.

jimthompson5910 (jim_thompson5910):

oh so 8.08 is wrong?

OpenStudy (anonymous):

actually it does here, 7.94

jimthompson5910 (jim_thompson5910):

hmm maybe they're using 58 days

jimthompson5910 (jim_thompson5910):

yeah 58, dang

OpenStudy (anonymous):

thats okay, now I know how to solve those. thanks!

jimthompson5910 (jim_thompson5910):

I = Prt I = 616.97(0.081)*(58/365) I = 7.9411645479452 I = 7.94